# TagAlign: Improving Vision-Language Alignment with Multi-Tag Classification

### v3

@@ 1 Introduction
Figure 1: (a)Illustration of the effect of various tag supervisions (e.g., object and attribute) for open-vocabulary semantic segmentation. The vanilla CLIP struggles in localizing an attribute-specified object. When introducing object supervision (as depicted in the third column), CLIP can focus on the more accurate region of text-specified object (e.g., cat). In addition, adding attribute supervision brings CLIP a stronger understanding of visual attribute-related concepts (as depicted in the fourth column). (b)Performance Comparison on various semantic segmentation benchmarks. TagAlign outperforms existing methods by a large margin on all these benchmarks. Best viewed in color.
Key aspect in the process of learning vision-language (V&L) encoders, e.g., CLIP [28], is to effectively extract semantically aligned information from both visual and linguistic data. A well-aligned vision-language encoder stands out for its remarkable versatility, which includes unprecedented quality in text-to-image generation, as well as the improvement in multi-modality understanding tasks. However, existing attempts usually face the problem of coarse alignment, e.g., the vision encoder struggles in localizing an attribute-specified object (as depicted in Fig. 1(a), making it hard to achieve precise localization of arbitrary concepts.
@@
In this paper, we propose an embarrassingly simple approach to improve the alignment between image and text features without the need for additional data formats beyond image-text pairs. Specifically, given an image and its paired text, we automatically parse objects (e.g., cat) and attributes (e.g., brown) from the description, which are highly likely to exist in the image. This parsing pipeline is implemented by a large language model, which is high-end and fully automatic, ensuring quality and scalability. By using these parsed semantics as supervision signals, we can complement the commonly used image-text contrastive loss with a multi-tag classification loss, enabling the model to better capture the visual semantic concepts referred to in the texts. Through that, our method becomes capable of accurately localizing text-specified objects (e.g., attribute-specified object), fulfilling more precise alignment between image and text. Extensive experimental results as shown in Fig. 1(b) on a broad suite of semantic segmentation datasets (10 datasets) substantiate the average 5.2% improvement of our framework over existing alternatives. Moreover, it is mentioned that we only use the CC12M dataset with the parsed text as the training set, indicating that our method is both efficient and effective.
Besides remarkable performance on open-vocabulary semantic segmentation tasks, TaskAlign also enjoys the high flexibility to incorporate attribute information. Concretely, the visualization results, as shown in Fig. 1(a), show that the inclusion of attribute supervision enables CLIP to accurately localize the attribute-specified object (‘brown cat’). This finding indicates that different types of tag supervision have the potential to enhance the ability of cross-modal alignment in different aspects of CLIP.

### v4

@@ 1 Introduction
Figure 1: Illustration of the effect of various tag supervisions (e.g., object and attribute) for open-vocabulary semantic segmentation. The vanilla CLIP struggles in localizing an attribute-specified object. When introducing object supervision (as depicted in the third column), CLIP can focus on the more accurate region of text-specified object (e.g., cat). In addition, adding attribute supervision brings CLIP a stronger understanding of visual attribute-related concepts (as depicted in the fourth column). Best viewed in color.
Figure 2: Performance Comparison on various semantic segmentation benchmarks. TagAlign outperforms existing methods by a large margin on all these benchmarks. Best viewed in color.
Key aspect in the process of learning vision-language (V&L) encoders, e.g., CLIP [28], is to effectively extract semantically aligned information from both visual and linguistic data. A well-aligned vision-language encoder stands out for its remarkable versatility, which includes unprecedented quality in text-to-image generation, as well as the improvement in multi-modality understanding tasks. However, existing attempts usually face the problem of coarse alignment, e.g., the vision encoder struggles in localizing an attribute-specified object (as depicted in Fig. 1), making it hard to achieve precise localization of arbitrary concepts.
@@
In this paper, we propose an embarrassingly simple approach to improve the alignment between image and text features without the need for additional data formats beyond image-text pairs. Specifically, given an image and its paired text, we automatically parse objects (e.g., cat) and attributes (e.g., brown) from the description, which are highly likely to exist in the image. This parsing pipeline is implemented by a large language model, which is high-end and fully automatic, ensuring quality and scalability. By using these parsed semantics as supervision signals, we can complement the commonly used image-text contrastive loss with a multi-tag classification loss, enabling the model to better capture the visual semantic concepts referred to in the texts. Through that, our method becomes capable of accurately localizing text-specified objects (e.g., attribute-specified object), fulfilling more precise alignment between image and text. Extensive experimental results as shown in Fig. 2 on a broad suite of semantic segmentation datasets (10 datasets) substantiate the average 5.2% improvement of our framework over existing alternatives. Moreover, it is mentioned that we only use the CC12M dataset with the parsed text as the training set, indicating that our method is both efficient and effective.
Besides remarkable performance on open-vocabulary semantic segmentation tasks, TaskAlign also enjoys the high flexibility to incorporate attribute information. Concretely, the visualization results, as shown in Fig. 1, show that the inclusion of attribute supervision enables CLIP to accurately localize the attribute-specified object (‘brown cat’). This finding indicates that different types of tag supervision have the potential to enhance the ability of cross-modal alignment in different aspects of CLIP.